Mechanics

The densities of graphite and diamond at 298K are 2.25 and 3.31gcm^-3 respectively. If the standard free energy difference is equal to 1895Jmol^-1, the pressure at which graphite will be transformed into diamond at 298K is
a. 9.92 * 10^7 Pa b. 9.92*10^6 Pa c. 9.92 * 10^5 Pa d. 9.9*10^8 pa

Regina Phalange
7 Years agoGrade

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
Free energy = useful work done∆G = p∆v1895= P(v2 - v1)1895= P(12/2.25 - 12/3.31) × 10^-6(m) × . 10^-3(kg) .......v=m/d1895= P(5.33 - 3.62) × 10^-91895= P × 1.71 × 10^-9 P= 1895/1.71×10^-9 P= 1108 × 10^9 P= 1.1 × 10^6